# Quantum Optimization Benchmark Library The Intractable Decathlon

### v1

@@ Abstract
Through recent progress in hardware development, quantum computers have advanced to the point where benchmarking of (heuristic) quantum algorithms at scale is within reach. Particularly in combinatorial optimization–where most algorithms are heuristics–it is key to empirically analyze their performance on hardware and track progress towards quantum advantage. To this extent, we present ten optimization problem classes that are difficult for existing classical algorithms and can (mostly) be linked to practically-relevant applications, with the goal to enable systematic, fair, and comparable benchmarks for quantum optimization methods. Further, we introduce the Quantum Optimization Benchmark Library (QOBLIB) [193] where the problem instances and solution track records can be found. The individual properties of the problem classes vary in terms of objective and variable type, coefficient ranges, and density. Crucially, they all become challenging for established classical methods already at system sizes ranging from less than $100$ to, at most, an order of 100 000100000100\,000100 000 decision variables, allowing to approach them with today’s quantum computers. We reference the results from state-of-the-art solvers for instances from all problem classes and demonstrate exemplary baseline results obtained with quantum solvers for selected problems. The baseline results illustrate a standardized form to present benchmarking solutions, which has been designed to ensure comparability of the used methods, reproducibility of the respective results, and trackability of algorithmic and hardware improvements over time. We encourage the optimization community to explore the performance of available classical or quantum algorithms and hardware platforms with the benchmarking problem instances presented in this work toward demonstrating quantum advantage in optimization.
@@ 1 Introduction
Optimization is subject of extensive research due to its importance in science and industry. Applications include experiment design, protein folding, logistics, and finance, where better algorithms can improve performance, accuracy, and cost efficiency. Optimization algorithms fall into three categories: provably exact algorithms, provably approximate algorithms with a priori performance guarantees, and heuristic algorithms without such guarantees, although posterior performance bounds may sometimes be available. Complexity theory classifies optimization problems as “easy” or “difficult”, determining what types of algorithms may exist. Many relevant optimization problems are NP-hard, i.e., the computational resources required to find provably optimal solutions grow exponentially with the problem size. We do not expect the existence of provably exact efficient algorithms for such problems in general [1]. For some of them, the possible performance of provably approximate algorithms is limited by known inapproximability bounds cf. [96, 128, 129]. Here, inapproximability bounds imply that solving a problem better than the given bound is NP-hard, i.e., as hard as solving the problem to optimality. In practice, it is often sufficient to find a good solution instead of a provably optimal one. Thus, heuristics are a viable approach for many problems of interest. Since heuristics lack a priori guarantees, rigorous benchmarking is crucial to assess the corresponding performance and resource requirements. To enable a fair comparison of algorithms and hardware platforms and to guarantee reproducibility of reported results, a suitable set of problem instances and metrics needs to be defined. This also allows tracking progress in algorithm and hardware improvements over time.
Quantum computing offers a novel toolbox to approach optimization problems–see [1] for an in-depth discussion of quantum optimization algorithms, their challenges, and potential. With recent advances in quantum hardware [126], systematic benchmarking of quantum optimization algorithms and comparison with classical solvers have become feasible. This paper proposes ten classes of combinatorial optimization problems, see Table 1, and corresponding problem instances, which can be found in an open-source repository Quantum Optimization Benchmark Library (QOBLIB) [193]. These problem classes were selected because they include instances that are difficult for state-of-the-art classical solvers with only a hundred to a hundred thousand variables, have diverse characteristics, and many of them have connections to industrial applications. This makes them suitable candidates for benchmarking and exploring the potential of quantum advantage in optimization. Further, for each class, we provide instances of varying complexity, ranging from those already feasible for today’s quantum hardware to those challenging for state-of-the-art classical solvers.
@@
Notably, we promote model-independent benchmarking, which allows for maximal flexibility in choosing models, algorithms, and hardware platforms. A prerequisite for our goal, i.e., the establishment of a fair and systematic benchmarking effort to advance toward demonstrating quantum advantage in optimization. Given the large number of possibilities, exhaustively comparing algorithms is an enormous challenge that requires the scientific community’s collective effort.
@@ 2 Related Work
Most benchmarks in quantum computing focus on the performance of individual components of the quantum computing stack and their integration [190, 2, 149]: Randomized benchmarking focus on gate quality [167]. The layer fidelity represents a quality metric of layers of two-qubit gates [168]. Circuit Layer Operations Per Second (CLOPS) measures the speed of a quantum computer [180, 242]. Speed and quality of compiling quantum circuits from high-level abstract representations to machine-level instructions have been investigated and reported in Benchpress [180]. In addition, application-driven benchmarks have been proposed as full-stack system benchmarks. They usually focus on problems that can be solved to optimality classically and track the progress of quantum computers to achieve a known optimal solution [83, 156, 142, 218, 209, 79]. While this allows tracking progress of hardware and algorithms, it is not sufficient to demonstrate quantum advantage.
This work presents a selected set of optimization problem instances that are difficult for state-of-the-art algorithms and corresponding benchmarking metrics that enable a fair and reliable comparison. See [163, 16, 140] for further discussion on the importance of appropriately chosen benchmarking metrics. The presented set of problem instances enables investigating potential advantages of novel algorithms and computational platforms. Furthermore, it facilitates tracking the progress towards potential quantum advantage in optimization. This is comparable to the curated set of many-body quantum systems that has been presented in [243] but for optimization.
@@ 3 Methods
In this section, we introduce our methodology to quantum optimization benchmarking. More specifically, we discuss what general problem types are considered, different ways to model them, and metrics to measure the performance of a specific optimization workflow.
@@ 3.1 Problem Types
From the theoretical side, both problem types are closely related, as we can solve optimization problems via a sequence of feasibility problems by adding a constraint for a certain objective value $k$, i.e., $X_{k}=\{x\in X\mid f(x)\leq k\}$, and asking whether $X_{k}=\emptyset$. The value of $k$ can then be iteratively adjusted, for example, using binary search, until the optimal value is found. We also can make the problem unconstrained, for instance, by removing the constraint and instead minimizing $\bar{f}(x)=f(x)+M\cdot\max\{0,g(x)-b\}$, $X=\{x\in\mathbb{Z}^{n}\mid l\leq x\leq u\}$ for a suitably large $M$, see Section 3.2 for a more detailed discussion.
@@ 3.2 Modeling
The choice of problem formulation, even in the same model class, can have a tremendous impact on the solvability of a problem [7]. While the global optimum is typically the same for different problem formulations [^1] Variations may occur, for example, in relaxed, e.g., linear or semi-definite programming formulations., one might be better suited for a particular problem instance or algorithm than another. The development of novel hardware and algorithmic devices to solve Quadratic Unconstrained Binary Optimization (QUBO) problems led to a renewed interest in formulations of problems as QUBO [154, 85]. While all Integer Linear Programs (ILP), or Integer Quadratic Programs (IQP) can be arbitrarily approximated by some QUBO, naive reformulations will often contain a large number of densely connected variables and will require outstanding precision for the coefficients to accurately represent the original structure [8]. For example, ILP can explicitly represent different constraints or variables that can be difficult to encode into a QUBO since many constraints can only be incorporated implicitly into QUBO via penalty terms that vanish for feasible solutions. In addition, to realize a QUBO representation, integer variables must be decomposed into binary ones–a useful library to handle this conversion can be found at [194]. Direct translation from a general ILP to a QUBO requires making decisions on hyperparameters that define the encoding and penalty functions [85] and can impact the quality of a solution depending on the considered algorithm. Furthermore, comparing the results received from multiple formulations of a problem requires careful interpretation. The objective values of two QUBOs representing the same problem using different penalty factors cannot be directly compared. Therefore, solutions must be mapped back to the original problem to understand and compare their performance. Higher-level modeling languages exist that provide additional structure to accommodate a problem’s original semantics more naturally, and multiple tools can automate reformulating from a high-level representation into, for instance, ILP or QUBO; see, e.g., [134, 248, 236]. Thus, the fact that a problem can be cast into a certain form does not necessarily mean that, in practice, it is a good idea to solve it in that form. Given the inherent formulation differences, we focus on model-independent benchmarking such that one may find the best formulation and corresponding algorithm for a given problem instance. Another type of reformulation allows mapping Higher-Order Unconstrained Binary Optimization (HUBOs), which may either be solved directly [184] or converted to QUBOs [4, 115]. To this extent, higher-order binary terms are broken into multiple quadratic terms by introducing additional binary variables. Depending on the number and degree of the higher-order terms, this can significantly increase the number of binary variables, potentially leading to impractical problem formulations. Some quantum optimization algorithms can handle higher-order terms natively, giving them a potential advantage over those algorithms (quantum and classical) that cannot. Notably, many quantum optimization algorithms are designed for QUBOs. A common approach is, thus, to start with a higher-level problem formulation and reformulate it as QUBO. Tutorials on QUBO reformulations can be found, e.g., in [85, 249].
@@ 3.3.2 Computational Resources
To compare different methods and algorithms, it is crucial to transparently report what type of computational resources have been used and for how long. Thus, software versions and hardware specifications, such as what type of processing units were used, how much memory was available, etc., must be reported, as discussed in more detail in the following.
First, the total runtime (s), i.e., the total wall-clock time for running the considered algorithm using the reported hardware, excluding idling times, e.g., queuing for execution on a (quantum) computer while the remaining processing units are waiting. For stochastic algorithms that are repeated multiple times, the average total runtime over all repetitions should be reported. If the algorithm is parallelized and uses multiple processing units simultaneously, this should be noted in the hardware description but not translated into a single-core runtime. The definition of the total runtime, excluding idling times, can be easily measured and represents the runtime most relevant for practical applications. In certain cases, it might also be useful to report the time-to-solution (TTS), i.e., the time from algorithm start until the solution that is returned was found.
Second, to better understand how different hardware is utilized, reported runtimes should be split into the execution time on CPUs, GPUs, quantum computers, and alternative hardware, such as, for instance, Field Programmable Gate Arrays (FPGAs) or Ising machines. How to accurately define, measure, and report the runtime on a quantum computer is discussed below in Section 3.3.3.
There are many additional details of interest, such as the time spent in different stages of the workflow, such as pre-solving, pre-processing (circuit preparation, etc.), and post-processing, or the intermediate problem sizes after pre-solving. While it is going too far to provide such details when submitting a benchmarking result, they could be shared in an accompanying reference such as a paper or repository.
@@ 3.3.3 Measuring Quantum Computing Runtimes
In the following, we define quantum computing runtime and how to measure it on the example of the IBM Quantum Platform [116, 162]; see Figure 1. A quantum computer consists of multiple stages, including a runtime environment, such as the Qiskit Runtime, the control electronics, and the actual quantum chip. The payload sent to a quantum computer needs to pass these stages, and even though some steps are performed using classical computing, such as, e.g., compiling a quantum circuit to microwave pulses, we attribute them to the runtime on the quantum computer. Thus, we define the runtime of a single run of a quantum (sub-) routine as the total time it takes to receive the payload, prepare it, run the circuit on the QPU, and extract and return samples from the system by measuring the qubits. We would like to highlight that the number of samples drawn from a quantum circuit corresponds to a hyperparameter that has to be chosen as part of the considered quantum algorithm. Furthermore, we assume that circuits are provided using a topology and gate set that is natively supported by the quantum computer. Thus, while we include the time it takes to prepare such a circuit for execution, e.g., to microwave pulses, we explicitly exclude the time it takes to map a given circuit to the instruction set supported by a particular quantum computer. This additional transpilation time should be considered during classical pre-processing. Similarly, we attribute time spent for error mitigation to the quantum computer, if it happens as part of the runtime environment, and to the classical processing units, if it is handled via post-processing.
@@ 4 Problem Classes & Instances
This section presents ten problem classes and corresponding instances that are non-trivial to solve with state-of-the-art methods–for most of them already at relatively small scales. Furthermore, the respective problem classes are mostly related to practical relevant problems. Such problems are hard to find for two reasons: First, unsolved problems of industrial relevance are rarely published. Second, problems of practical relevance that cannot be solved are usually discarded and replaced by artificial simplifications of the original problem. This leads to a selection bias and may be one reason why a benchmarking framework that includes truly difficult (for reasonable system sizes) and relevant (ideally in practical settings) optimization problems that enable a meaningful and fair comparison of quantum and classical algorithms has not been established until now.
Problem instances can be loosely classified into three categories, which are not always mutually exclusive: Real-world instances arise from real-world problems, and are therefore crucial for demonstrating a practical quantum advantage. Since there is already a demand for a solution, such instances have often been modeled (and possibly simplified) such that they can be tackled by existing solvers. The availability of real-world instances to the scientific community is often limited, even more for unsolved problems as discussed above. Random instances are based on randomly generated data, which may be motivated by real-world problems or completely artificial. Depending on the generation method, such instances often do not capture the structural peculiarities of a real-world setting, and may be easy or hard to solve. Crafted instances are most likely the best way to construct artificial examples that are hard to solve. However, there is a risk that the instances are of no practical relevance, as they often represent borderline pathological cases; see, e.g., [114, 229]. A quantum advantage for random or crafted instances is interesting, as it may help to identify real-world problems with similar structure. The problem instances presented here were mostly crafted to fit a range of criteria we identified as crucial for a useful optimization benchmarking framework, while they are still mostly related to real-world problems.
@@
Market Split F / L d 78 $48$ $1$ no yes d 70 $\sim 5\cdot 10^{6}$
LABS Q / Q s 81 4 $1$ yes yes s 820 $\sim 4\cdot 10^{4}$
Birkhoff L / L s 240 $10^{4}$ $3$ yes no d 3,480 $\sim 3\cdot 10^{10}$
Steiner L / L s 423,360 $3$ $9$ no yes - - -
Sports F / L s 8,608 $2$ $>10$ no no s $11,791$ $\sim 4.5\cdot 10^{3}$
Portfolio Q / L s 690 $\sim 3\cdot 10^{4}$ $2$ yes yes d 690 $\sim 2\cdot 10^{9}$
Independent Set L / L s 500 $1$ $1$ yes yes d 500 2
Network L / L s 1,211 $10^{6}$ $5$ yes no s 46,330 $\sim 2.5\cdot 10^{19}$
Routing L / L s - - $<10$ yes no s - -
Topology L / Q s 2,176 $2$ $4-7$ yes no s - -
@@
Table 2: Problem classes overview - details on illustrative MIP and QUBO formulations, as applicable. Type denotes the objective/constraint type as linear (L) or quadratic (Q) for either if the problem corresponds to an optimization problem. For feasibility problems, the objective type is indicated as feasibility (F). Dense denotes whether the model is dense (d), defined as $|A|\geq\frac{1}{4}nm$, where $A\in\mathbb{R}^{m\times n}$ is the constraint matrix of a MIP and $n,m$ denote the number of variables and constraints, respectively, or sparse (s), otherwise. For QUBO models, density is defined as $|Q|>\frac{1}{8}n^{2}$, where we assume $Q$ is an upper-triangular cost matrix (due to symmatry, and thus, the reduced pre-factor of $\frac{1}{8}$ instead of $\frac{1}{4}$ used for MIP). Again, $n$ denotes the number of variables. Portfolio has sparse constraints matrix but a dense quadratic objective, which we indicate by $s^{*}$. #Vars denotes the approximate number of decision variables of the models corresponding to instances where the problems can become difficult for classical methods to solve to proven optimality in MIP formulation. Hence, the number of decision variables in the QUBO column denotes the size of the QUBO formulation where solvers struggle to solve the corresponding MIP. This is to highlight how certain problems become significantly larger when represented as QUBO instead of a MIP. Coeffs is the approximate coefficient range, i.e., the maximum range of all coefficients in the models. Again, we can have a significant discrepancy between MIP and QUBO representations for certain classes. Constr Number of different constraint types in the MIP formulation ($\neq$ total number of constraints). This indicates the complexity of modeling the problems. Feas denotes whether it is trivial to find or construct a (initial) feasible solution to the problem or not. For QUBO, all solutions are feasible, and thus, we drop the column. Bin indicates whether all variables in the MIP problem are naturally binary or not. This is always true for QUBO and, thus, not reported. If this is not the case, we see a discrepancy between the number of variables in the MIP and the QUBO models. For LABS, we have natural binary variables, however, we receive quartic terms in the objective from adding squared quadratic equality constraints as penalties that need to be decomposed into quadratic terms by adding binary variables. QUBO instances requiring more than $1.6\cdot 10^{5}$ variables have not been generated due to computational restrictions. We indicate this in the respective columns (-). Further, no values are given for Routing problems since we were able to find the optimal solution for all tested instances.
@@
  • Portfolio Optimization and Vehicle Routing problems have obvious practical applications. Companies face diverse variations of these problems depending on industry and operation execution. Hence, there are indefinitely many variations of these problems depending on individual constraints and/or objectives, resulting in problem instances of varying difficulty. For this reason, it is difficult to define a small set of problem instances to benchmark and compare established approaches, which are representative for these problem classes. In addition, many established solvers from commercial software vendors are not accessible for open-source benchmarks. Despite the practical relevance of these problem classes, no well-established benchmark set exists for testing the progress of quantum algorithms. We aim to fill this gap by providing a selection of relevant problem instances. However, it should be noted that all given Vehicle Routing problem instances can be solved to optimality with existing methods. This is a shortcoming that we aim to address in the future.

### v2

# Quantum Optimization Benchmarking Library The Intractable Decathlon

@@ Abstract
Through recent progress in hardware development, quantum computers have advanced to the point where benchmarking of (heuristic) quantum algorithms at scale is within reach. Particularly in combinatorial optimization–where most algorithms are heuristics–it is key to empirically analyze their performance on hardware and track progress towards quantum advantage. To this extent, we present ten optimization problem classes that are difficult for existing classical algorithms and can (mostly) be linked to practically relevant applications, with the goal to enable systematic, fair, and comparable benchmarks for quantum optimization methods. Further, we introduce the Quantum Optimization Benchmarking Library [202] where the problem instances and solution track records can be found. The individual properties of the problem classes vary in terms of objective and variable type, coefficient ranges, and density. Crucially, they all become challenging for established classical methods already at system sizes ranging from less than $100$ to, at most, an order of 100 000100\,000 decision variables, allowing to approach them with today’s quantum computers. We reference the results from state-of-the-art solvers for instances from all problem classes and demonstrate exemplary baseline results obtained with quantum solvers for selected problems. The baseline results illustrate a standardized form to present benchmarking solutions, which has been designed to ensure comparability of the used methods, reproducibility of the respective results, and trackability of algorithmic and hardware improvements over time. We encourage the optimization community to explore the performance of available classical or quantum algorithms and hardware platforms with the benchmarking problem instances presented in this work toward demonstrating quantum advantage in optimization.
@@ 1 Introduction
Optimization is subject to extensive research due to its importance in science and industry. Applications include experiment design, protein folding, logistics, and finance, where better algorithms can improve performance, accuracy, and cost efficiency. Optimization algorithms fall into three categories: provably exact algorithms, provably approximate algorithms with a priori performance guarantees, and heuristic algorithms without such guarantees, although posterior performance bounds may sometimes be available. Complexity theory classifies optimization problems as “easy” or “difficult”, determining what types of algorithms may exist. Many relevant optimization problems are NP-hard, i.e., the computational resources required to find provably optimal solutions grow exponentially with the problem size. We do not expect the existence of provably exact efficient algorithms for such problems in general [1]. For some of them, the possible performance of provably approximate algorithms is limited by known inapproximability bounds cf. [103, 134, 136]. Here, inapproximability bounds imply that solving a problem better than the given bound is NP-hard, i.e., as hard as solving the problem to optimality. In practice, it is often sufficient to find a good solution instead of a provably optimal one. Thus, classic heuristics and ML-based solvers are viable approaches for many problems of interest. Since heuristics lack a priori guarantees, rigorous benchmarking is crucial to assess the corresponding performance and resource requirements. To enable a fair comparison of algorithms and hardware platforms and to guarantee reproducibility of reported results, a suitable set of problem instances and metrics needs to be defined. This also allows tracking progress in algorithm and hardware improvements over time.
Quantum computing offers a novel toolbox to approach optimization problems–see [1] for an in-depth discussion of quantum optimization algorithms, their challenges, and potential. With recent advances in quantum hardware [132], systematic benchmarking of quantum optimization algorithms and comparison with classical solvers have become feasible. This paper proposes ten classes of combinatorial optimization problems, see Table 1, and corresponding problem instances, which can be found in an open-source repository Quantum Optimization Benchmark Library [202]. These problem classes were selected because they include instances that are difficult for state-of-the-art classical solvers with only a hundred to a hundred thousand variables, have diverse characteristics, and many of them have connections to industrial applications. This makes them suitable candidates for benchmarking and exploring the potential of quantum advantage in optimization. Further, for each class, we provide instances of varying complexity, ranging from those already feasible for today’s quantum hardware to those challenging for state-of-the-art classical solvers.
@@
Notably, we promote model-independent benchmarking, which allows for maximal flexibility in choosing models, algorithms, and hardware platforms. A prerequisite for our goal is the establishment of a fair and systematic benchmarking effort to advance toward demonstrating quantum advantage in optimization. Given the large number of possibilities, exhaustively comparing algorithms is an enormous challenge that requires the scientific community’s collective effort.
@@ 2 Related Work
Most benchmarks in quantum computing focus on the performance of individual components of the quantum computing stack and their integration [199, 2, 159]: Randomized benchmarking focuses on gate quality [176]. The layer fidelity represents a quality metric of layers of two-qubit gates [177]. Circuit Layer Operations Per Second (CLOPS) measures the speed of a quantum computer [189, 252]. Speed and quality of compiling quantum circuits from high-level abstract representations to machine-level instructions have been investigated and reported in Benchpress [189]. In addition, application-driven benchmarks have been proposed as full-stack system benchmarks. They usually focus on problems that can be solved to optimality classically and track the progress of quantum computers to achieve a known optimal solution [88, 165, 150, 227, 218, 84]. While this allows tracking the progress of hardware and algorithms, it is not sufficient to demonstrate quantum advantage.
This work presents a selected set of optimization problem instances that are difficult for state-of-the-art algorithms and corresponding benchmarking metrics that enable a fair and reliable comparison. See [172, 16, 148] for further discussion on the importance of appropriately chosen benchmarking metrics. The presented set of problem instances enables investigating potential advantages of novel algorithms and computational platforms. Furthermore, it facilitates tracking the progress towards potential quantum advantage in optimization. This is comparable to the curated set of many-body quantum systems presented in [253]–but for optimization.
@@ 3 Methods
In this section, we introduce our methodology for quantum optimization benchmarking. More specifically, we discuss what general problem types are considered, different ways to model them, and metrics to measure the performance of a specific optimization workflow.
@@ 3.1 Problem Types
From the theoretical side, both problem types are closely related, as we can solve optimization problems via a sequence of feasibility problems by adding a constraint for a certain objective value $k$, i.e., $X_{k}=\{x\in X\mid f(x)\leq k\}$, and asking whether $X_{k}=\emptyset$. The value of $k$ can then be iteratively adjusted, for example, using binary search, until the optimal value is found. We can also make the problem unconstrained, for instance, by removing the constraint and instead minimizing $\bar{f}(x)=f(x)+M\cdot\max\{0,g(x)-b\}$, $X=\{x\in\mathbb{Z}^{n}\mid l\leq x\leq u\}$ for a suitably large $M$, see Section 3.2 for a more detailed discussion.
@@ 3.2 Modeling
The choice of problem formulation, even in the same model class, can have a tremendous impact on the solvability of a problem [7]. While the global optimum is typically the same for different problem formulations [^1] Variations may occur, for example, in relaxed, e.g., linear or semi-definite programming formulations., one might be better suited for a particular problem instance or algorithm than another. The development of novel hardware and algorithmic devices to solve Quadratic Unconstrained Binary Optimization (QUBO) problems led to a renewed interest in formulations of problems as QUBO [163, 91]. While all Integer Linear Programs (ILP), or Integer Quadratic Programs (IQP) can be arbitrarily approximated by some QUBO, naive reformulations will often contain a large number of densely connected variables and will require outstanding precision for the coefficients to accurately represent the original structure [8]. For example, ILP can explicitly represent different constraints or variables that can be difficult to encode into a QUBO, since many constraints can only be incorporated implicitly into QUBO via penalty terms that vanish for feasible solutions. In addition, to realize a QUBO representation, integer variables must be decomposed into binary ones–a useful library to handle this conversion can be found at [203]. Direct translation from a general ILP to a QUBO requires making decisions on hyperparameters that define the encoding and penalty functions [91] and can impact the quality of a solution depending on the considered algorithm. Furthermore, comparing the results received from multiple formulations of a problem requires careful interpretation. The objective values of two QUBOs representing the same problem using different penalty factors cannot be directly compared. Therefore, solutions must be mapped back to the original problem to understand and compare their performance. Higher-level modeling languages exist that provide additional structure to accommodate a problem’s original semantics more naturally, and multiple tools can automate reformulating from a high-level representation into, for instance, ILP or QUBO; see, e.g., [141, 258, 245]. Thus, the fact that a problem can be cast into a certain form does not necessarily mean that, in practice, it is a good idea to solve it in that form. Given the inherent formulation differences, we focus on model-independent benchmarking such that one may find the best formulation and corresponding algorithm for a given problem instance. Another type of reformulation allows mapping Higher-Order Unconstrained Binary Optimization (HUBOs), which may either be solved directly [193] or converted to QUBOs [4, 121]. To this extent, higher-order binary terms are broken into multiple quadratic terms by introducing additional binary variables. Depending on the number and degree of the higher-order terms, this can significantly increase the number of binary variables, potentially leading to impractical problem formulations. Some quantum optimization algorithms can handle higher-order terms natively, giving them a potential advantage over those algorithms (quantum and classical) that cannot. Notably, many quantum optimization algorithms are designed for QUBOs. A common approach is, thus, to start with a higher-level problem formulation and reformulate it as QUBO. Tutorials on QUBO reformulations can be found, e.g., in [91, 259].
@@ 3.3.1 Solution Quality
Notably, one may also want to track metrics that present combined information about solution quality and the time it takes to arrive at a particular quality. One example thereof is the primal integral [20], which tracks the primal bound towards the optimal (or best known) solution throughout the iterations executed by a solver. This approach has been found to be suitable for comparing the performance of solvers that do not necessarily return optimal solutions. This includes machine learning-based optimization solvers [146, 31, 25, 158]–trained on a large number of known instances–which can, in certain cases, predict good solutions quickly.
@@ 3.3.2 Computational Resources
To compare different methods and algorithms, it is crucial to transparently report the type of computational resources used and the duration for which they were utilized. Thus, software versions and hardware specifications, such as the type of processing units used and the amount of memory available, must be reported, as discussed in more detail below.
First, the total runtime (s), i.e., the total wall-clock time for running the considered algorithm using the reported hardware, excluding idling times, e.g., queuing for execution on a (quantum) computer while the remaining processing units are waiting. For stochastic algorithms that are repeated multiple times, the average total runtime over all repetitions should be reported. If the algorithm is parallelized and uses multiple processing units simultaneously, this should be noted in the hardware description but not translated into a single-core runtime. The definition of the total runtime, excluding idling times, can be easily measured and represents the runtime most relevant for practical applications. In certain cases, it may also be useful to report the time-to-solution (TTS), i.e., the time from the algorithm’s start until the returned solution was found.
Second, to better understand how different hardware is utilized, reported runtimes should be split into the execution time on CPUs, GPUs, quantum computers, and alternative hardware, such as, for instance, Field Programmable Gate Arrays (FPGAs) or Ising machines. How to accurately define, measure, and report runtime on a quantum computer is discussed in Section 3.3.3 below.
There are many additional details of interest, such as the intermediate problem sizes after pre-solving or the time and resources spent in different stages of the workflow: pre-solving, pre-processing (circuit preparation, etc.), training of machine learning-based solvers, post-processing, etc. While it is difficult to capture all of these details in a submission table (see Table 10), we strongly encourage benchmark submitters to provide such details in an accompanying reference, such as a paper or repository.
@@ 3.3.3 Measuring Quantum Computing Runtimes
In the following, we define quantum computing runtime and how to measure it on the example of the IBM Quantum Platform [122, 171]; see Figure 1. A quantum computer consists of multiple stages, including a runtime environment, such as the Qiskit Runtime, the control electronics, and the actual quantum chip. The payload sent to a quantum computer needs to pass these stages, and even though some steps are performed using classical computing, such as, e.g., compiling a quantum circuit to microwave pulses, we attribute them to the runtime on the quantum computer. Thus, we define the runtime of a single run of a quantum (sub-) routine as the total time it takes to receive the payload, prepare it, run the circuit on the QPU, and extract and return samples from the system by measuring the qubits. We would like to highlight that the number of samples drawn from a quantum circuit corresponds to a hyperparameter that has to be chosen as part of the considered quantum algorithm. Furthermore, we assume that circuits are provided using a topology and gate set that is natively supported by the quantum computer. Thus, while we include the time it takes to prepare such a circuit for execution, e.g., to microwave pulses, we explicitly exclude the time it takes to map a given circuit to the instruction set supported by a particular quantum computer. This additional transpilation time should be considered during classical pre-processing. Similarly, we attribute time spent on error mitigation to the quantum computer, if it happens as part of the runtime environment, and to the classical processing units, if it is handled via post-processing.
@@ 4 Problem Classes & Instances
This section presents ten problem classes and corresponding instances that are non-trivial to solve with state-of-the-art methods–for most of them already at relatively small scales. Furthermore, the respective problem classes are mostly related to practically relevant problems. Such problems are hard to find for two reasons: First, unsolved problems of industrial relevance are rarely published. Second, problems of practical relevance that cannot be solved are usually discarded and replaced by artificial simplifications of the original problem. This leads to a selection bias and may be one reason why a benchmarking framework that includes truly difficult (for reasonable system sizes) and relevant (ideally in practical settings) optimization problems that enable a meaningful and fair comparison of quantum and classical algorithms has not been established until now.
Problem instances can be loosely classified into three categories, which are not always mutually exclusive: Real-world instances arise from real-world problems, and are therefore crucial for demonstrating a practical quantum advantage. Since there is already a demand for a solution, such instances have often been modeled (and possibly simplified) such that they can be tackled by existing solvers. The availability of real-world instances to the scientific community is often limited, even more so for unsolved problems as discussed above. Random instances are based on randomly generated data, which may be motivated by real-world problems or completely artificial. Depending on the generation method, such instances often do not capture the structural peculiarities of a real-world setting, and may be easy or hard to solve. Crafted instances are most likely the best way to construct artificial examples that are hard to solve. However, there is a risk that the instances are of no practical relevance, as they often represent borderline pathological cases; see, e.g., [120, 238]. A quantum advantage for random or crafted instances is interesting, as it may help to identify real-world problems with similar structure. The problem instances presented here were mostly crafted to fit a range of criteria we identified as crucial for a useful optimization benchmarking framework, while they are still mostly related to real-world problems.
@@
Market Split F / L d 78 $48$ 11 no yes d 70 $\sim 5\cdot 10^{6}$
LABS Q / Q s 81 4 11 yes yes s 820 $\sim 4\cdot 10^{4}$
Birkhoff L / L s 240 $10^{4}$ 33 yes no d 3,480 $\sim 3\cdot 10^{10}$
Steiner L / L s 423,360 33 99 no yes - - -
Sports F / L s 8,608 22 >10>10 no no s $11,791$ $\sim 4.5\cdot 10^{3}$
Portfolio Q / L s 690 $\sim 3\cdot 10^{4}$ 22 yes yes d 690 $\sim 2\cdot 10^{9}$
Independent Set L / L s 500 11 11 yes yes d 500 2
Network L / L s 1,211 $10^{6}$ 55 yes no s 46,330 $\sim 2.5\cdot 10^{19}$
Routing L / L s - - <10<10 yes no s - -
Topology L / Q s 2,176 22 474-7 yes no s - -
@@
Table 2: Problem classes overview - details on illustrative MIP and QUBO formulations, as applicable. Type denotes the objective/constraint type as linear (L) or quadratic (Q) for either if the problem corresponds to an optimization problem. For feasibility problems, the objective type is indicated as feasibility (F). Dense denotes whether the model is dense (d), defined as $|A|\geq\frac{1}{4}nm$, where $A\in\mathbb{R}^{m\times n}$ is the constraint matrix of a MIP and $n,m$ denote the number of variables and constraints, respectively, or sparse (s), otherwise. For QUBO models, density is defined as $|Q|>\frac{1}{8}n^{2}$, where we assume $Q$ is an upper-triangular cost matrix (due to symmetry, and thus, the reduced pre-factor of $\frac{1}{8}$ instead of $\frac{1}{4}$ used for MIP). Again, $n$ denotes the number of variables. Portfolio has a sparse constraints matrix but a dense quadratic objective, which we indicate by $s^{*}$. #Vars denotes the approximate number of decision variables of the models corresponding to instances where the problems can become difficult for classical methods to solve to proven optimality in MIP formulation. Hence, the number of decision variables in the QUBO column denotes the size of the QUBO formulation, where solvers struggle to solve the corresponding MIP. This is to highlight how certain problems become significantly larger when represented as QUBO instead of an MIP. Coeffs is the approximate coefficient range, i.e., the maximum range of all coefficients in the models. Again, we can have a significant discrepancy between MIP and QUBO representations for certain classes. Constr Number of different constraint types in the MIP formulation ($\neq$ total number of constraints). This indicates the complexity of modeling the problems. Feas denotes whether it is trivial to find or construct a (initial) feasible solution to the problem or not. For QUBO, all solutions are feasible, and thus, we drop the column. Bin indicates whether all variables in the MIP problem are naturally binary or not. This is always true for QUBO and, thus, not reported. If this is not the case, we see a discrepancy between the number of variables in the MIP and the QUBO models. For LABS, we have natural binary variables; however, we receive quartic terms in the objective from adding squared quadratic equality constraints as penalties that need to be decomposed into quadratic terms by adding binary variables. QUBO instances requiring more than $1.6\cdot 10^{5}$ variables have not been generated due to computational restrictions. We indicate this in the respective columns (-). Further, no values are given for Routing problems since we were able to find the optimal solution for all tested instances.
@@
  • Portfolio Optimization and Vehicle Routing problems have obvious practical applications. Companies face diverse variations of these problems depending on industry and operational execution. Hence, there are indefinitely many variations of these problems depending on individual constraints and/or objectives, resulting in problem instances of varying difficulty. For this reason, it is difficult to define a small set of problem instances to benchmark and compare established approaches, which are representative of these problem classes. In addition, many established solvers from commercial software vendors are not accessible for open-source benchmarks. Despite the practical relevance of these problem classes, no well-established benchmark set exists for testing the progress of quantum algorithms. We aim to fill this gap by providing a selection of relevant problem instances. However, it should be noted that all given Vehicle Routing problem instances can be solved to optimality with existing methods. This is a shortcoming that we aim to address in the future.